HA and AH represent two 2-digit numbers. If HA - AH =18, what is the value of the expression H - A?
Answer by Edwin McCravy(20054)

(10H + A) - (10A + H) = 18

    10H + A - 10A - H = 18

              9H - 9A = 18

Divide through by 9

                H - A = 2

--------------------------------------------

Note that there is no way to find H and A.  In fact
there are seven possibilities:

HA = 31 and AH = 13 or
HA = 42 and AH = 24 or
HA = 53 and AH = 35 or
HA = 64 and AH = 46 or
HA = 75 and AH = 57 or
HA = 86 and AH = 68 or
HA = 97 and AH = 79

But in each of these 7 cases, H - A = 2
 
Edwin